# From Large to Super-Tiny: End-to-End Optimization for Cost-Efficient LLMs

### v1

@@ Appendix G Detailed Analysis for Experimental Results for KD
As demonstrated in Table 7, it is evident that balancing FKL and RKL is quite crucial for our task. The process of function selection by the model resembles a classification task, whereas the generation of parameter names and values for function calls is more akin to a generation task. The parameter $\mu$ in AKL is precisely used to adjust the effects between classification and generation. Experimental results demonstrate that at $\mu\leq 0.7$ and $\mu=1.0$ (equivalent to FKL), the performance with AKL is similar to or worse than RFT. Based on our findings where the average probability of the top-1 token from the teacher model exceeds 99%, we draw the conclusion that setting a lower $\mu$ value causes the effect of AKL to excessively shift more towards RKL. As described in the AKL paper, RKL prioritizes learning the tail of the probability distribution, which is detrimental for rapid convergence in scenarios without requiring diversity. However, RKL can bring considerable benefits for the function classification part. Overall, AKL demonstrates greater potential compared to RFT, achieving some performance improvements with an appropriately chosen $\mu$ value.
@@ Appendix H Inference Details
Appendix I Prune Detail
@@ I.1 Offline One-shot Pruning
Importance Metrics for Structural Pruning
(1) Depth Pruning: we propose Layer Importance (LI) metric based on the similarity between consecutive layers’ activation. Formally, we define the importance of the $i$-th layer as:
$LI_{i}=1-E_{X,t}\frac{X_{i,t}^{T}X_{i+1,t}}{\|X_{i,t}\|_{2}\|X_{i+1,t}\|_{2}}$ (11)
where $X_{i}$ denotes the activation matrix of layer $i$, with $X_{i,t}$ representing the $t$-th column (embedding vector) of $X_{i}$. After we collecte all layer activations $\{X_{i}\}_{i=1}^{L}$ through a single forward pass on calibration data, then remove the $N$ least important layers, where $N$ is determined by the target pruning ratio; (2) Width Pruning: We categorize width pruning along three dimensions by channel dependencies: Embedding dimension, Head dimension and Intermediate dimension(as as illustrated in ). For compact LLMs (e.g., Qwen2.5-0.5B with head size 14), we exclude head dimension pruning due to its demonstrated sensitivity[37], we use an activation-based importance metric Channel Importance(CI) to do Channel selection as:
@@
where $CI^{i}_{\text{embedding}}$ denotes the importance score of the $i$-th embedding dimension, computed from LayerNorm outputs of attention and FFN blocks, while $CI^{i}_{\text{inter}}$ represents the importance of the $i$-th intermediate dimension, calculated from the $i$-th row of FFN’s gate and up projection matrices $W_{ffn1}^{i}$. The importance scores are first aggregated across Batch and Sequence dimensions through summation $\sum_{B,S}$, then normalized using L2 normalization within each subgroup to obtain global channel importance.
Figure 8: Pruning dependency in Qwen
Best Width Pruning Model Configuration
To maintain hardware-efficient matrix multiplication during inference, we align pruning granularity to 64 for Embedding dimension and 256 for intermediate dimension, yielding multiple width-pruned model configurations as shown in Table 8. We compute channel importance scores $CI^{(i)}_{\text{embedding}}$ and $CI^{(j)}_{\text{inter}}$ on a calibration dataset comprising 1024 randomly sampled sequences from WikiText-2. For each configuration, we prune entire channel groups with the lowest CI scores and evaluate the resulting one-shot pruned models through WikiText-2 perplexity. Table 8 presents the optimal configurations for width pruning at rates of 20% and 30%, which are identified as Qwen2.5-0.4B-Width-2 and Qwen2.5-0.35B-Width-3, respectively.
@@ I.2 Recovery Training in Pruning
General Instruction Fine-tuning Dataset And Traing Hyperparameters
We employ the Infinity-Instruct-7M[106] dataset and adopt AKL distillation[29] described in Section 3.2 for general instruction recovery training. We utilize Qwen2.5-7B-Instruct[11] as the teacher model, following Minitron’s methodology[67] to perform teacher correction on the Infinity-Instruct-7M dataset. We employ the Adam optimizer (weight decay=1e-4) with a cosine learning rate schedule decaying from 2e-5 to 2e-6. We conduct training for 2 epochs, utilizing a batch size of 32, and each batch is packed to accommodate sequences of up to 7168 tokens.
Table 9: Performance of different pruned model compared to the unpruned model. We report 5-shot performance on mmlu and cmmlu, 4-shot for gsm8k, zero-shot for ARC-Challenge, HellaSwag. * indicates that the baseline Qwen-2.5-0.5B utilizes the same instruct fine-tuning dataset for correction.
@@
Evaluation on Knowledge & Reasoning Tasks
Table 9 presents the performance of pruned models on downstream tasks, including MMLU and CMMLU for knowledge assessment, as well as ARC-Challenge, HellaSwag and Winogrande for commonsense reasoning evaluation. Our experiments demonstrate distinct recovery patterns between width and depth pruning on general tasks. At 20% pruning rate, width pruning achieves 90.5% performance retention compared to the unpruned baseline, while depth pruning retains only 78.4%. When increasing pruning intensity to 30%, width pruning maintains 83.8% of unpruned performance, whereas depth pruning suffers catastrophic performance degradation-notably on GSM8K, where accuracy drops from 45.11 to merely 5.61.
@@
Our investigation reveals a nonlinear relationship between pruning intensity and model accuracy in domain-specific tasks. Table 10 demonstrates that both depth-wise and width-wise pruning with 20% pruning rate incur negligible AR degradation (<0.5%). However, when increasing the pruning rate to 30%, we observe significant performance deterioration (>1.5% AR drop). These findings suggest that for small-scale models operating on domain-specific tasks, pruning rates exceeding 20% may lead to substantial accuracy compromises.
Table 10: Results of different pruning ratios and pruning strategies on domain specific task.
@@
  • $B$: batch size
  • $H$: hidden size
  • $S$: sequence length (average number of prompt tokens)
  • $S_{\text{uncached}}$: average number of tokens in prefill phase not hit by prefill cache (requires recomputing $kv_{\text{cache}}$)
  • attn_heads: number of attention heads
  • kv_heads: number of separate key/value heads in GQA
  • head_size: dimension of each head
  • inter_size: intermediate dimension of FFN layer
  • vocab_size: vocabulary size
  • $L$: number of model layers
Following [122], we consider only the floating-point operations (FLOPs) from matrix multiplications (GEMMs). For the Qwen model:
@@
    • Query/Key/Value transformations (GQA):
[…]
    • Attention matrix computation and attention over values:
[…]
    • Post-attention linear projection:
[…]
@@
    • Up/Gate/Down projections (each):
[…]
@@
[…]
@@
Our experiments workload consists of prompts with an average length of 1792 tokens. By leveraging prefix caching to exploit shared prompt prefixes during inference, we reduce the average number of uncached tokens in the prefill phase to 128, while generating an average of 9 tokens per request. For FLOPs calculations, we account for $S_{\text{uncached}}=128$ tokens in the prefill phase and $S_{\text{uncached}}=1$ token per decoding step, with detailed results presented in Table 2.

### v2

@@ Appendix G Detailed Analysis for Experimental Results for KD
As demonstrated in Table 7, it is clear that balancing FKL and RKL is crucial for our task. The function selection process by the model resembles a classification task, whereas generating parameter names and values for function calls is more akin to a generation task. The parameter $\mu$ in AKL is precisely used to adjust the effects between classification and generation. Experimental results demonstrate that at $\mu\leq 0.7$ and $\mu=1.0$ (equivalent to FKL), the performance with AKL is similar to or worse than RFT. Based on our findings where the average probability of the top-1 token from the teacher model exceeds 99%, we conclude that setting a lower $\mu$ value causes the effect of AKL to excessively shift more towards RKL. As described in the AKL paper, RKL prioritizes learning the tail of the probability distribution, which can impede rapid convergence in scenarios without requiring diversity. However, RKL can bring considerable benefits for the function classification part. Overall, AKL demonstrates greater potential compared to RFT, achieving performance improvements with an appropriately chosen $\mu$ value.
@@ Appendix H Inference Details
Appendix I Pruning Details
@@ I.1 Offline One-shot Pruning
Importance as Metric for Structured Pruning
For Depth Pruning, we adopt Layer Importance (LI) as the metric, which is based on the similarity between consecutive layers’ activation. Formally, we define the importance of the $i$-th layer as:
$LI_{i}=1-\mathbb{E}_{X,t}\frac{X_{i,t}^{T}X_{i+1,t}}{\|X_{i,t}\|_{2}\|X_{i+1,t}\|_{2}}$ (11)
where $X_{i}$ denotes the activation matrix of the $i$-th layer and $X_{i,t}$ denotes the $t$-th column (embedding vector) of $X_{i}$. After all layer activations $\{X_{i}\}_{i=1}^{L}$ through a single forward pass on calibration data are collected, the $N$ least important layers are removed, where $N$ is determined by the target pruning ratio. For Width Pruning, we classify width pruning into three dimensions based on channel dependencies: Embedding dimension, Head dimension and Intermediate dimension (as illustrated in Figure 8). For compact LLMs (e.g., Qwen2.5-0.5B with head size of 14), we exclude head dimension pruning due to its demonstrated sensitivity[37]. Instead, we employ an activation-based importance metric, Channel Importance(CI), for channel selection as follows:
@@
where $CI^{i}_{\text{embedding}}$ denotes the importance score of the $i$-th embedding dimension, computed from the LayerNorm outputs of attention and FFN blocks, while $CI^{i}_{\text{inter}}$ denotes the importance of the $i$-th intermediate dimension, calculated from the $i$-th row of FFN’s gate and up projection matrices $W_{ffn1}^{i}$. The importance scores are first aggregated across batch and sequence dimensions through summation $\sum_{b,s}$, then normalized using L2 normalization within each subgroup to obtain global channel importance.
Figure 8: Pruning Dependency in the Qwen2.5
Optimal Width Pruning Configuration
To ensure hardware-efficient matrix multiplication during inference, we set the pruning granularity to 64 for embedding dimension and 256 for the intermediate dimension. This approach results in multiple width-pruned model configurations, as show in Table 8. We calculate channel importance scores $CI^{(i)}_{\text{embedding}}$ and $CI^{(j)}_{\text{inter}}$ using a calibration dataset consisting of 1024 randomly sampled sequences from WikiText-2. For each configuration, we prune entire channel groups with the lowest CI scores and evaluate the resulting one-shot pruned models through WikiText-2 perplexity. Table 8 presents the optimal configurations for width pruning at rates of 20% and 30%, identified as Qwen2.5-0.4B-Width-2 and Qwen2.5-0.35B-Width-3, respectively.
@@ I.2 Recovery Training in Pruning
Recovery Training Dataset and Hyperparameters
We use the Infinity-Instruct-7M[83] dataset and apply AKL distillation[29] described in Section 3.2 for general instruction recovery training. Qwen2.5-7B-Instruct[11] serves as the teacher model, and we follow Minitron’s methodology[71] to perform teacher correction on the Infinity-Instruct-7M dataset. The Adam optimizer is used with a weight decay of 1e-4 in conjunction with a cosine learning rate schedule, decaying from 2e-5 to 2e-6. We conduct training over 2 epochs, utilizing a batch size of 32, with each batch packed to accommodate sequences of up to 7168 tokens.
Table 9: Performance comparison of Qwen2.5 models after pruning. We report 5-shot performance on mmlu and cmmlu, 4-shot for gsm8k, zero-shot for ARC-Challenge, HellaSwag. * denotes that the baseline Qwen-2.5-0.5B utilizes the same instruct fine-tuning dataset for correction.
@@
Performance across Knowledge Assessment & Reasoning Tasks
Table 9 summarizes the performance of pruned models across various downstream tasks, including MMLU and CMMLU for knowledge assessment, alongside ARC-Challenge, HellaSwag and Winogrande for commonsense reasoning evaluation. Our experiments reveal distinct recovery patterns between width and depth pruning across general tasks. At a 20% pruning rate, width pruning achieves 90.5% of the performance relative to the unpruned baseline, while depth pruning retains only 78.4%. When increasing the pruning intensity to 30%, width pruning preserves 83.8% of the original performance. In contrast, depth pruning suffers catastrophic performance degradation, particularly on the GSM8K task, where accuracy drops from 45.11 to a merely 5.61.
@@
Our investigation reveals a nonlinear relationship between pruning intensity and model accuracy in domain-specific tasks. As shown in Table 10, both depth-wise and width-wise pruning at a 20% pruning rate result in a negligible <0.5% AR reduction. However, increasing the pruning rate to 30% leads to notable performance deterioration (>1.5% AR drop). These findings suggest that for small-scale models on domain-specific tasks, pruning rates exceeding 20% may significantly compromise accuracy.
Table 10: Results with different pruning strategies on domain specific tasks.
@@
  • $b$: the batch size
  • $h$: the hidden size
  • $s_{\text{t}}$: the total average number of prompt tokens
  • $s_{\text{u}}$: the average number of uncached tokens in prefill phase
  • $n_{\text{a}}$: the number of attention heads
  • $n_{\text{kv}}$: the number of separate key/value heads in GQA
  • $d_{\text{h}}$: the dimension of each head
  • $d_{\text{i}}$: the intermediate dimension of FFN layer
  • $v$: the vocabulary size
  • $l$: the number of model layers
According to [84], we calculate the floating-point operations (FLOPs) of the Qwen2.5 models, only considering matrix multiplications (GEMMs) without norms and activations:
@@
    • Query/Key/Value transformations (GQA): $2bs_{\text{u}}hd_{\text{h}}(n_{\text{a}}+2n_{\text{kv}})$
    • Attention matrix computation and attention over values: $2bs_{\text{t}}s_{\text{u}}n_{\text{a}}d_{\text{h}}$
    • Post-attention linear projection: $2bs_{\text{u}}hn_{\text{a}}d_{\text{h}}$
@@
    • Up/Gate/Down projections (each): $2bs_{\text{u}}hd_{\text{i}}$
  • LM head operations: $2bhv$
@@
$\text{FLOPs}_{\ \text{total}}$ $=l\Bigg{[}4bs_{\text{u}}hd_{\text{h}}\left(n_{\text{a}}+n_{\text{kv}}\right)+2bs_{\text{u}}s_{\text{t}}d_{\text{h}}n_{\text{a}}+6bs_{\text{u}}hd_{\text{i}}\Bigg{]}+2bhv$
@@
We conduct experiments on our datasets, in which prompts have an average length of 1792 tokens. By leveraging prefix caching to exploit shared prompt prefixes during inference, we reduce the average number of uncached tokens in the prefill phase to 128 and generate an average of 9 tokens per request. We calculate FLOPs under $s_{\text{u}}=128$ tokens in the prefill phase and under $s_{\text{u}}=1$ token for each step in the decoding phase, with detailed results presented in Table 2.
Appendix K Extensive Attempts on Generalization Enhancement via RL
Table 11: Full results during the RL exploration
Model AR
AR
(Out-of-Domain)
Crowdsourced Labelers 77.85% -
BERT-based System 81.43% -
Qwen2.5-72B (ICL) 94.74% 94.63%
Qwen2.5-7B (RFT) 95.29% 90.86%
Qwen2.5-3B (RFT) 95.30% 91.02%
Qwen2.5-0.5B (RFT) 94.66% 90.62%
Qwen2.5-0.5B (KD) 94.66% 90.62%
Qwen2.5-7B (RL) 95.36% 94.80%
Qwen2.5-3B (RL) 94.68% 92.94%
Qwen2.5-0.5B (RL) 90.76% 85.21%
Qwen2.5-0.5B (KD*) 88.49% 84.14%
Qwen2.5-0.5B (KD*+RL) 95.55% 91.68%
While attempting to apply reinforcement learning (RL), we conduct extensive experiments, with results shown in Appendix K. Initially, we find that directly training a 7B model with RL yields excellent results, surpassing the LLM Prototype and RFT-trained models. Subsequently, we attempt to train smaller models (3B and 0.5B) and discover that RL is ineffective on the 0.5B model, because the model tends to generate shorter or even meaningless Chain-of-Thought (CoT), leading to degraded output diversity as well as that without CoT. To address the issue above, we explore another approach, transferring out-of-domain capabilities into the 0.5B model via Knowledge Distillation (KD). However, we observe that the model struggles with format errors due to underfitting. Ultimately, after applying RL training on this model, we notice an impressive improvement on model performance, with both an absolute 1% AR increase on in-domain/out-of-domain datasets compared to the 0.5B RFT-trained model.